OpenStudy (anonymous):

Factor Completely 2sinx cos y- 3secx tan y- sin x tan y + 6cos y sec x

OpenStudy (anonymous):

rewrite it and i can help

OpenStudy (anonymous):

its hard for me to help in that form

OpenStudy (anonymous):

Factor: 2sinxcosy - 3secxtany - sinxtany + 6cosysecx Does this help? I'm not sure how else to write it.

OpenStudy (aum):

2sinx cos y - 3secx tany - sin x tan y + 6cos y sec x = 2sinx cos y - sin x tan y + 6cos y sec x - 3secx tany = sinx(2cosy - tany) + 3secx(2cosy - tany) = (sinx + 3secx)(2cosy - tany)

OpenStudy (anonymous):

Thank You!

OpenStudy (aum):

You are welcome.